# Synthetic Datasets for Machine Learning on Spatio-Temporal Graphs using PDEs

### v1

@@ A.4 Creation of Epidemiological Real-World Datasets
    The Brazilian COVID-19 dataset has 27 nodes and 1093 time-steps spanning 2019-2022 after concatenation and linear interpolation for daily resolution and is publicly accessible. 333sisaps.saude.gov.br/painelsaps/atendimento
@@ Appendix B Details on: Applications of Synthetic Data
The exact hyperparameters can be found in the implementation on GitHub in the regarding parameter files such as github.com/github-usr-ano/Temporal_Graph_Data_PDEs/blob/main/ml/mp_pde/mp_pde.yml for the MP-PDE model. The parameter files for other models can be found in their respective directories. The hyperparameters were chosen as optimal for each model experimentally. We trained all models for each task with the Adam optimizer until convergence (early stopping).
@@ Appendix C Details on Epidemiological Benchmarking
Figure 10: The RMSE over all samples from the test dataset. The less saturated colors represent the standard deviation for different initializations.
@@
Figure 11: Plots of the RMSE of the presented models during the stability experiments for each predicted time-step. The models were trained on the synthetic dataset, and evaluated on a synthetic dataset with: a) additive Gaussian noise, b) dropout noise. The less saturated colors represent the standard deviation for different initializations.
@@
Figure 12: Plots of the RMSE of the presented models during the denoising experiments for each predicted time-step. The models were trained and evaluated on a synthetic dataset which. The models were trained to predict based on context data, with a) additive Gaussian noise, b) dropout noise. The less saturated colors represent the standard deviation for different initializations.
@@ E.1 Motivation
All involved people are listed as authors of this paper. The authors are all employees of Fraunhofer Heinrich Hertz Institut, Berlin, Germany.
@@ E.3 Collection Process
All involved people are listed as authors of this paper and receive compensation as they are employees of Fraunhofer Heinrich Hertz Institut. To list them here again: Jost Arndt, Utku Isil, Michael Detzel, Samek Wojciech, and Jackie Ma.

### v2

@@ A.4 Creation of Epidemiological Real-World Datasets
    The Brazilian COVID-19 dataset has 27 nodes and 1093 time-steps spanning 2019-2022 after concatenation and linear interpolation for daily resolution and is publicly accessible. 555sisaps.saude.gov.br/painelsaps/atendimento
@@ Appendix B Details on: Applications of Synthetic Data
The exact hyperparameters can be found in the implementation on GitHub in the regarding parameter files, exemplary under /ml/mp_pde/mp_pde.yml for the MP-PDE model. The parameter files for other models can be found in their respective directories. The hyperparameters were chosen as optimal for each model experimentally. We trained all models for each task with the Adam optimizer until convergence (early stopping).
@@ Appendix C Details on Epidemiological Benchmarking
Figure 10: The RMSE over all samples from the validation dataset. The less saturated colors represent the standard deviation for different initializations.
@@
Figure 11: Plots of the RMSE of the presented models during the stability experiments for each predicted time-step on the validation dataset. The models were trained on the synthetic dataset, and evaluated on a synthetic dataset with: a) additive Gaussian noise, b) dropout noise. The less saturated colors represent the standard deviation for different initializations.
@@
Figure 12: Plots of the RMSE of the presented models during the denoising experiments for each predicted time-step on the validation dataset. The models were trained and evaluated on the synthetic dataset. The models were trained to predict based on context data, with a) additive Gaussian noise, b) dropout noise. The less saturated colors represent the standard deviation for different initializations.
@@ E.1 Motivation
All involved people are listed as authors of this paper. The authors are all employees of Fraunhofer Heinrich Hertz Institute, Berlin, Germany.
@@ E.3 Collection Process
All involved people are listed as authors of this paper and receive compensation as they are employees of Fraunhofer Heinrich Hertz Institute. To list them here again: Jost Arndt, Utku Isil, Michael Detzel, Wojciech Samek, and Jackie Ma.